OpenStudy (anonymous):

f(x)=x^3−5x^2+5x−3 when is the function concave up and when is it concave down

OpenStudy (anonymous):

not you u dumb bimbo did u give that guy nudes or not?

OpenStudy (anonymous):

To know that, you need to look at the second derivative of the function. If it is positive, then its concave up, if negative concave down. \[\frac{d^2f}{dx^2}=6x-10\]The second derivative is zero at x=10/6 so f(x) is concave down if x<10/6 and concave up if x>10/6. To understand why the positive is concave up, and the negative concave down, you need to look at the second derivative as the rate of change of the slope of the line, if the second derivative is positive, it means that the slope is increasing, therefore it is concave up, if it is negative, the other way around.

OpenStudy (anonymous):

thank you